Suppose the production cost is R1895, labour cost is and equipment renting fee is R 250, determine the number of devices produced



1
Expert's answer
2022-03-02T16:25:34-0500

Production Cost = PC = 1895

Labor Cost = LC = 250

Rental Cost = RC = 250


Say the no. of devices = N

Unit Production cost (BoM Cost) = "\frac{(PC + RC + LC)}{N}" 


(BoM --> Bill of Materials)


total cost for N no. of devices = 1895 + 250 + 250 = 2395


Unit cost = "\frac{Total cost}{No. of Devices}"


Therefore, Unit Cost = "\frac{2395}{N}"


Hence no. of Device = N = "\frac{Total Cost}{Unit Cost}"



I have provided the solution using the available information in the problem. However, if the unit cost is provided, then, the absolute no. of devices may be calculated. Unit cost of the product is missing in the problem. Therefore, I have put the missing information as a variable UC. Once UC is put in the formula, absolute no. of devices may be calculated.